OpenStudy (anonymous):

HELPP ASAP how do you graph this? f(x) =|x-5|

OpenStudy (anonymous):

The lines on both sides of x-5 are 'absolute marks' this means whatever value comes from x-5 is always positive.

OpenStudy (anonymous):

So, for example 2-5 = -3 but \[\left| 2-5 \right| = 3\]

OpenStudy (anonymous):

using that information, do you think you can make a graph of that function?

OpenStudy (anonymous):

of?

OpenStudy (anonymous):

x-5?

OpenStudy (anonymous):

Well, ok let's do that

OpenStudy (anonymous):

sketch a graph of x-5 please.

OpenStudy (anonymous):



OpenStudy (anonymous):

alright, seems about right assuming you meant a straight line

OpenStudy (anonymous):